Bernice Invested $120,000. She put some in an account that pays 4% interest per year and the remainder she invested in an account that pays 7% per year. If she earned $7,800 in one year, how much was invested at each rate? Write an equation.
Answer by stanbon(75887)

Bernice Invested $120,000. She put some in an account that pays 4% interest per year and the remainder she invested in an account that pays 7% per year. If she earned $7,800 in one year, how much was invested at each rate? Write an equation.
--------
Equation:
interest + interest = interest
0.04x + 0.07(120,000-x) = 7800
----
4x + 7*120000 - 7x = 780000
------------------
-3x = -60000
x = $20,000 (amt. invested at 4%
120,000-x = $100,000 (amt. invested at 7%)
==================
Cheers,
Stan H.
==================